| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  194  
 | 
Fri Jul  2 03:10:03 2021 | 
IK, KK | Runs | drop of intensity | We see that intensity stabalized at 9E8 which 
is smaller then before. 
 
We neticed this at 2:50, keep observing. | 
   | 
| 
  193  
 | 
Fri Jul  2 02:03:36 2021 | 
The remote Canadian beavers (Iris, Guy) | Analysis | Summary of spectra 0:00-1:00 | Same as before, for 0-1h on July 2nd
 
 
 | 
    | 
| 
  192  
 | 
Fri Jul  2 01:43:56 2021 | 
IK, KK | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000360
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0001759
 | 
   | 
| 
  191  
 | 
Fri Jul  2 01:39:51 2021 | 
IK, KK | Runs | drop of intensity | 
 
We see a drop in beam intensity from 7E9 
to 2E8 between 01:30 and 01:43 | 
   | 
| 
  190  
 | 
Fri Jul  2 01:34:03 2021 | 
The remote Canadian beavers (Iris, Guy) | Analysis | Summary of spectra 22:00-23:00 | The last hour before the gas bottle change 
is summarized here. 
 
The first pic shows the 410 MHz spectra taken | 
      | 
| 
  189  
 | 
Fri Jul  2 01:31:35 2021 | 
The remote Canadian beavers (Iris, Guy) | Analysis | First spectra after bottle change | The first good spectra after the gas bottle 
change is from 23:38:25 (on the 410 MHz Analyzer) 
 | 
     | 
| 
  188  
 | 
Fri Jul  2 00:08:22 2021 | 
IK, Kanika | DAQ | NTCAP status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000338
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0001651
 | 
   | 
| 
  187  
 | 
Thu Jul  1 23:54:43 2021 | 
Iris | General | Less beam after gas bottle change? | No, everything okay! Must have had a very 
good earlier run to compare.
 
 | 
    | 
| 
  186  
 | 
Thu Jul  1 23:46:17 2021 | 
Iris | General | Less beam after gas bottle change? | Looks like there is a bit less beam now in 
the ring. I just looked at one spectra from 
before the gas bottle change and after... | 
    | 
| 
  185  
 | 
Thu Jul  1 23:38:55 2021 | 
IK, KK | Runs | Change of gas bottle | The gas bottle has been changed and the beam 
is back. | 
   | 
| 
  184  
 | 
Thu Jul  1 23:07:47 2021 | 
IK, KK | Runs | source | 
 
The change of the source happens now | 
   | 
| 
  183  
 | 
Thu Jul  1 21:35:32 2021 | 
Shahab | Runs | some pics after 15dB amplification on 410 |   | 
     | 
| 
  182  
 | 
Thu Jul  1 21:35:15 2021 | 
Shahab | DAQ | increase factor 4 in power for 410  | increase factor 4 in power for 410 
 
 
 | 
   | 
| 
  181  
 | 
Thu Jul  1 21:15:01 2021 | 
Andrew Nic | DAQ | NTCAP Status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000299
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0001462
 | 
   | 
| 
  180  
 | 
Thu Jul  1 20:48:58 2021 | 
Shahab Andrew Nic | DAQ | Acquisition started | 15dB amplifier on 410MHz  | 
   | 
| 
  179  
 | 
Thu Jul  1 20:35:21 2021 | 
Andrew Nic Shahab | DAQ | Acquisition stopped  | Checking if signal can be further amplified 
 
 | 
   | 
| 
  178  
 | 
Thu Jul  1 19:34:11 2021 | 
Shahab | DAQ | increase factor 4 in power for 410  | increase factor 4 in power for 410 
 
 
 | 
     | 
| 
  177  
 | 
Thu Jul  1 19:24:59 2021 | 
Nic Andrew Shahab | Runs | Start of data taking again | 410 MHz resolution and power optimised
 
 
 | 
   | 
| 
  176  
 | 
Thu Jul  1 18:56:49 2021 | 
Shahab | Runs | Short interruption in recording  | DAQs and NTCAP | 
   | 
| 
  175  
 | 
Thu Jul  1 18:54:44 2021 | 
Andrew, Nic | General | NTCap Status | The latest recorded tdms files:
 
 p:\E143\sc\SC_2021-06-30_23-27-45\0000267
 
 p:\E143\iq\IQ_2021-06-30_23-27-34\0001305
 | 
   |